if (2)(3)(7) / (5)(6)(7) = x/y and (3)(4)(5) / (6)(7)(8) = ax/by, what is the value of a/b ?

A 25/16 B 25/29 C 4 D 49/16 E 1/49

Answer by Alan3354(69443)

if (2)(3)(7) / (5)(6)(7) = x/y and (3)(4)(5) / (6)(7)(8) = ax/by, what is the value of a/b ?
-------------------
(2)(3)(7) / (5)(6)(7) = x/y = 1/5
(3)(4)(5) / (6)(7)(8) = ax/by = 5/28
--------------
a/b = (5/28)/(1/5) = 25/28